OpenStudy (anonymous):

If I have a square, and one side is 4x-11, and the other is x+7 what is x?

OpenStudy (anonymous):

If it's a square, both sides are equal, therefore: 4x-11=x+7 then just solve for x

OpenStudy (anonymous):

I know, but I dont know how, thats why I asked.

OpenStudy (anonymous):

I think its 6.

OpenStudy (anonymous):

that's correct

OpenStudy (anonymous):

Move the lone x to the left and the -11 to the right. When you do that, they change sign, so: 4x-x=7+11 doing the operations: 3x=18 therefore x=18/3 or x=6

OpenStudy (anonymous):

\[4x-11=x+7\] \[4x-11+11=x+7+11\] \[4x=x+18\] \[4x-x=x-x+18\] \[x(4-1)=18\] \[3x=18\] \[x=\frac{18}{3}=6\]

OpenStudy (anonymous):

Thank you.